# Moser’s twist theorem revisited

### v2

@@ 4.2. Difference quotients
$\lambda(v_{1},v_{2},v_{3},v_{4}):=\frac{1}{2}(\lambda(v_{1},v_{3})+\lambda(v_{2},v_{4})).$
@@
$\Theta(v_{1},\ldots,v_{4}):=\|\theta(v_{1},\ldots,v_{4})\|.$
@@
$\frac{1}{(A_{\alpha}+1)^{6}}q_{s-2\gamma_{0}}\leq|i_{3}-i_{1}|\leq(A_{\alpha}+1)^{6}q_{s-2\gamma_{0}},$
$\lambda(v_{1},v_{2},v_{3},v_{4})\in[\|q_{s-2\gamma_{0}}\alpha\|,2^{4}\|q_{s-2\gamma_{0}}\alpha\|),$
where $\gamma_{0}$ is defined by (9).
@@
$\nabla^{1}(v_{1},v_{2}):=\frac{s(v_{2})-s(v_{1})}{\Theta(v_{1},v_{2})},$
@@
$\nabla^{1}(\kappa,r):=\sup_{(v_{1},v_{2})\in(\kappa,r)_{II}}|\nabla^{1}(v_{1},v_{2})|.$
$\nabla^{2}(\kappa,r,s):=\sup_{(v_{1},\ldots,v_{4})\in(\kappa,r,s)_{II}}|\nabla^{2}(v_{1},\ldots,v_{4})|.$
@@
(17) $E^{1}(j|v_{1},v_{2}):=\Theta(v_{1},v_{2})^{-1}\left(\frac{\Theta(F^{j}(v_{2}))}{\Theta(v_{2})}-\frac{\Theta(F^{j}(v_{2}))}{\Theta(v_{2})}\right).$
@@
Proof   Note that $\lambda$ is an invariant metric with respect to $F$. For each $j\in\mathbb{Z}$, we have
$\lambda(F^{j}(v_{i}))=\lambda(v_{i})\ (i=1,2),\quad\lambda(F^{j}(v_{1}),F^{j}(v_{2}))=\lambda(v_{1},v_{2}).$
@@
$\Theta(v)=\|x_{l+q_{\kappa}}-x_{l}\|<\|x_{l+q_{n_{0}}}-x_{l}\|\sim\frac{1}{q_{n_{0}}},\quad\mathrm{for}\ \kappa\geq n_{0}.$
@@
$\frac{\Theta(F(v_{2}))}{\Theta(v_{2})}-\frac{\Theta(F(v_{1}))}{\Theta(v_{1})}$
@@
$\frac{s(v_{2})-s(v_{1})}{\Theta(v_{1},v_{2})}=\nabla^{1}(v_{1},v_{2}).$
[…]
@@ 4.3. Growth of the distortions
$\tilde{\mathbb{K}}_{r}^{1}(\kappa,r)\lesssim(\frac{9}{10})^{\frac{\varepsilon}{3}r}.$
For consistency, we first verify Criterion 3 by using Lemma 4.3. Let us recall $\bar{N}:=\bar{N}(\kappa)$ is defined by (9). For a given $j$ with $0\leq j<q_{\bar{N}}$, it can be written as
@@
$\tilde{\mathbb{K}}^{0}_{\bar{N}}(\kappa)$ $\leq c\mathbb{K}^{0}_{\tilde{N}}(\kappa)+\sum_{r=0}^{\tilde{N}-1}a_{r}\mathbb{K}^{0}_{r}(\kappa)$
$\leq 2c\tilde{\mathbb{K}}^{1}_{\tilde{N}}(\kappa,\tilde{N})+2\sum_{r=0}^{\tilde{N}-1}a_{r}\tilde{\mathbb{K}}^{1}_{r}(\kappa,r).$
@@
$c\leq\frac{q_{\bar{N}}}{q_{\tilde{N}}}=\frac{q_{n_{\kappa}+2\gamma_{0}}}{q_{n_{\kappa}-1}}\sim 1.$
@@
$\tilde{\mathbb{K}}^{0}_{\bar{N}}(\kappa)$ $\lesssim\tilde{\mathbb{K}}^{1}_{\tilde{N}}(\kappa,\tilde{N})+\sum_{r=0}^{\tilde{N}-1}\tilde{\mathbb{K}}^{1}_{r}(\kappa,r),$
$\lesssim(\frac{9}{10})^{\frac{\varepsilon}{3}\tilde{N}}+\ldots+(\frac{9}{10})^{\frac{\varepsilon}{3}r}+\ldots+1,$
$\lesssim\sum_{r=0}^{+\infty}(\frac{9}{10})^{\frac{\varepsilon}{3}r},$
which implies $\tilde{\mathbb{K}}_{\bar{N}(\kappa)}^{0}(\kappa)\lesssim 1$.
@@
Denote
$K^{2}(v_{1},\ldots,v_{4}):=\frac{K^{1}(v_{3},v_{4})}{\Theta(v_{3},v_{4})}-\frac{K^{1}(v_{1},v_{2})}{\Theta(v_{1},v_{2})};$
$K^{2}(j|v_{1},\ldots,v_{4}):=\frac{K^{1}(j|v_{3},v_{4})}{\Theta(v_{3},v_{4})}-\frac{K^{1}(j|v_{1},v_{2})}{\Theta(v_{1},v_{2})}.$
It follows that
(22) $\Theta(v_{3},v_{4})K^{2}(j|v_{1},\ldots,v_{4})=K^{1}(j|v_{3},v_{4})-\frac{\Theta(v_{3},v_{4})}{\Theta(v_{1},v_{2})}K^{1}(j|v_{1},v_{2}).$
Let $v\in(\kappa)_{II}$ for $\kappa$ large enough. Note that if $v=v^{\prime}+v^{\prime\prime}$, then $\Theta(F^{j}(v))=\Theta(F^{j}(v^{\prime}))+\Theta(F^{j}(v^{\prime\prime}))$ for each $j\in\mathbb{Z}$. Moreover,
$\frac{\Theta(F^{j}(v))}{\Theta(v)}=\frac{\Theta(v^{\prime})}{\Theta(v)}\frac{\Theta(F^{j}(v^{\prime}))}{\Theta(v^{\prime})}+\frac{\Theta(v^{\prime\prime})}{\Theta(v)}\frac{\Theta(F^{j}(v^{\prime\prime}))}{\Theta(v^{\prime\prime})}.$
@@
Similarly, for any $r\leq\kappa$, we have
(24) $\nabla^{1}(\kappa,r)\leq\nabla^{1}(\kappa+1,r)\leq\nabla^{1}(\kappa+1,r+1),$
(25) $\nabla^{2}(\kappa,r,s)\leq\nabla^{2}(\kappa+1,r,s)\leq\nabla^{2}(\kappa+1,r+1,s).$
@@
    There holds $\tilde{\mathbb{K}}_{r}^{1}(\kappa,R)\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}$.
@@
    $K^{2}(j|v_{1},\ldots,v_{4})-\sum_{i=0}^{j-1}K^{2}(F^{i}(v_{1},\ldots,v_{4}))\lesssim\varepsilon e^{\frac{\varepsilon}{10}R}q_{r}q_{R}^{-1}\sum_{i=0}^{j-1}\frac{\Theta(F^{i}(v_{3},v_{4}))}{\Theta(v_{3},v_{4})}.$
@@
    (26) $e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}\leq e^{\frac{\varepsilon}{10}R}q_{r}q_{R}^{-1}\sim e^{\frac{\varepsilon}{10}R}q_{R}^{-\frac{\varepsilon}{2}}\leq(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
Proof   By (8), $q_{R}\geq\sqrt{2}^{R}$. For Item (3), a direct calculation shows
@@
For $(v_{1},v_{2})\in(\kappa,r)_{II}$ with $0\leq r\leq R$, denote by $J(v_{1},v_{2})$ the arc on $\mathbb{T}$ spanned by $v_{1}\cup v_{2}$, then for $0\leq j<q_{r}$, no point on $\mathbb{T}$ is contained in more than two of the arcs $\{F^{i}(J(v_{1},v_{2}))\}_{i=0}^{j}$. It implies
$\sum_{i=0}^{j}\Theta(F^{i}(v_{1}),F^{i}(v_{2}))\leq 2.$
@@
$=\sum_{i=0}^{j-1}\left(\ln\frac{\Theta(F^{i+1}(v_{2}))}{\Theta(F^{i}(v_{2}))}-\ln\frac{\Theta(F^{i+1}(v_{1}))}{\Theta(F^{i}(v_{1}))}\right).$
By (12), $\frac{\Theta(F^{i+1}(v))}{\Theta(F^{i}(v))}$ is bounded for each $v\in(\kappa)_{II}$ and $i\in\mathbb{Z}$. It follows from the mean value theorem and Lemma 4.2 that for $0\leq j<q_{r}$,
(27) $K^{1}(j|v_{1},v_{2})\leq 3\nabla^{1}(\kappa,r)\sum_{i=0}^{j}\Theta(F^{i}(v_{1}),F^{i}(v_{2}))\leq\frac{\varepsilon}{80A_{\alpha}}.$
Note that $r\leq R\leq n_{\kappa}$. It follows from (24) and (19) that
$\tilde{\mathbb{K}}_{r}^{1}(r,r)\leq\frac{\varepsilon}{80A_{\alpha}},\quad\tilde{\mathbb{K}}_{r}^{1}(\kappa,r)\leq\frac{\varepsilon}{80A_{\alpha}}.$
By Proposition 4.1 and (21), for $0<j<q_{R}$, we have
$K^{0}(j|v)\leq 2a_{R}\tilde{\mathbb{K}}_{R-1}^{1}(\kappa,R-1)+\ldots+2a_{l+1}\tilde{\mathbb{K}}_{l}^{1}(\kappa,l)+\ldots+2a_{1}\tilde{\mathbb{K}}_{0}^{1}(\kappa,0),$
which together with $1\leq a_{l}\leq A_{\alpha}$ implies
(28) $\tilde{\mathbb{K}}_{R}^{0}(\kappa)\leq\frac{\varepsilon R}{40}.$
By definition, there exists a chord $\bar{v}$ determined by $v_{1},v_{2}$ such that $\lambda(\bar{v})=\lambda(v_{1},v_{2})$ and $\bar{v}\in({\phi(R-2\gamma_{0})})_{II}$. It follows that $\lambda(\bar{v})\geq\frac{2}{q_{R}}$. Note that
$\tilde{\mathbb{K}}_{R}^{0}({\phi(R-2\gamma_{0})})\leq\tilde{\mathbb{K}}_{R}^{0}({\phi(R)})\leq\tilde{\mathbb{K}}_{R}^{0}(\kappa).$
It follows from a similar argument as Lemma 3.7 that
$-\frac{\varepsilon R}{40}\leq\ln\frac{\Theta(\bar{v})}{\lambda(\bar{v})}\leq\frac{\varepsilon R}{40}.$
@@
$K^{1}(j|v_{1},v_{2})\lesssim\varepsilon\sum_{i=0}^{q_{r}}\Theta(F^{i}(v_{1}),F^{i}(v_{2}))\leq\varepsilon e^{\frac{\varepsilon R}{40}}\sum_{i=0}^{q_{r}}\lambda(F^{i}(v_{1},v_{2}))\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}.$
@@
(29) $\tilde{\mathbb{K}}_{r}^{1}(R,R)\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1},\quad\tilde{\mathbb{K}}_{r}^{1}(\kappa,R)\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}.$

### v3

@@ 4.2. Difference quotients
$\lambda_{4}(v_{1},v_{2},v_{3},v_{4}):=\frac{1}{2}(\lambda(v_{1},v_{3})+\lambda(v_{2},v_{4})).$
@@
$\Theta_{4}(v_{1},\ldots,v_{4}):=\|\theta(v_{1},\ldots,v_{4})\|.$
@@
$\frac{1}{q_{s+2}}\leq\lambda_{4}(v_{1},v_{2},v_{3},v_{4})\leq\frac{1}{q_{s}}.$
@@
$\nabla^{1}(v_{1},v_{2}):=\frac{s(v_{2})-s(v_{1})}{\Theta_{2}(v_{1},v_{2})},$
@@
(23) $\nabla^{1}(\kappa,r):=\sup_{(v_{1},v_{2})\in(\kappa,r)_{II}}|\nabla^{1}(v_{1},v_{2})|.$
(24) $\nabla^{2}(\kappa,r,s):=\sup_{(v_{1},\ldots,v_{4})\in(\kappa,r,s)_{II}}|\nabla^{2}(v_{1},\ldots,v_{4})|.$
@@
(25) $E^{1}(j|v_{1},v_{2}):=\Theta_{2}(v_{1},v_{2})^{-1}\left(\frac{\Theta_{1}(F^{j}(v_{2}))}{\Theta_{1}(v_{2})}-\frac{\Theta_{1}(F^{j}(v_{2}))}{\Theta_{1}(v_{2})}\right).$
@@
Proof.
Note that $\lambda$ is an invariant metric with respect to $F$. For each $j\in\mathbb{Z}$, we have
$\lambda_{1}(F^{j}(v_{i}))=\lambda(v_{i})\ (i=1,2),\quad\lambda_{2}(F^{j}(v_{1}),F^{j}(v_{2}))=\lambda_{2}(v_{1},v_{2}).$
@@
$\Theta_{1}(v)=\|x_{l+q_{\kappa}}-x_{l}\|<\|x_{l+q_{n_{0}}}-x_{l}\|\sim\frac{1}{q_{n_{0}}},\quad\mathrm{for}\ \kappa\geq n_{0}.$
@@
$\frac{\Theta_{1}(F(v_{2}))}{\Theta_{1}(v_{2})}-\frac{\Theta_{1}(F(v_{1}))}{\Theta_{1}(v_{1})}$
@@
$\frac{s(v_{2})-s(v_{1})}{\Theta_{2}(v_{1},v_{2})}=\nabla^{1}(v_{1},v_{2}).$
This completes the proof of Lemma 4.2. ∎
@@ 4.3. Growth of the distortions
(29) $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,r)\lesssim(\frac{9}{10})^{\frac{\varepsilon}{2}r}.$
For consistency, we first verify Criterion 3 by using the conclusion of Lemma 4.3. Let us recall $\bar{N}:=\bar{N}(\kappa)$ is defined by (12). For a given $j$ with $0\leq j<q_{\bar{N}}$, it can be written as
@@
(31) $\begin{split}\widetilde{\mathbf{K}}^{0}_{\bar{N}}(\kappa)&\leq c\mathbb{K}^{0}_{\tilde{N}}(\kappa)+\sum_{r=0}^{\tilde{N}-1}a_{r}\mathbb{K}^{0}_{r}(\kappa)\\ &\leq 2c\widetilde{\mathbf{K}}^{1}_{\tilde{N}}(\kappa,\tilde{N})+2\sum_{r=0}^{\tilde{N}-1}a_{r}\widetilde{\mathbf{K}}^{1}_{r}(\kappa,r).\end{split}$
@@
(32) $c\leq\frac{q_{\bar{N}}}{q_{\tilde{N}}}=\frac{q_{n_{\kappa}+2\gamma_{0}}}{q_{n_{\kappa}-2\gamma_{0}}}\sim 1.$
@@
(33) $\begin{split}\widetilde{\mathbf{K}}^{0}_{\bar{N}}(\kappa)&\lesssim\widetilde{\mathbf{K}}^{1}_{\tilde{N}}(\kappa,\tilde{N})+\sum_{r=0}^{\tilde{N}-1}\widetilde{\mathbf{K}}^{1}_{r}(\kappa,r),\\ &\lesssim(\frac{9}{10})^{\frac{\varepsilon}{2}\tilde{N}}+\ldots+(\frac{9}{10})^{\frac{\varepsilon}{2}r}+\ldots+1,\\ &\lesssim\sum_{r=0}^{+\infty}(\frac{9}{10})^{\frac{\varepsilon}{2}r},\end{split}$
which implies $\widetilde{\mathbf{K}}_{\bar{N}(\kappa)}^{0}(\kappa)\lesssim 1$.
@@
To help the reader follow the argument, we first outline the main steps.
  • (1)
    Lemma 4.4 provides monotonicity properties of $\nabla^{1}(\kappa,r)$ with respect to $\kappa$ and $r$.
  • (2)
    The proof proceeds in two stages. Lemma 4.6 supplies three basic estimates that are needed to establish (29); none of them uses condition (28). Specifically,
    • Lemma 4.6(1) gives a rough upper bound for $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,r)$;
    • Lemma 4.6(2) is used to obtain (54) in the proof of Lemma 4.8; for a more compact presentation we include it as part of Lemma 4.6(2);
    • Lemma 4.6(3) follows from a direct computation and clarifies under which relation between $r$ and $R$ (roughly $r\ll R$) the rough bound of Lemma 4.6(1) actually yields exponential decay of $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)$.
  • (3)
    Lemma 4.6 shows that when $r\ll R$, one already obtains exponential decay of $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)$. Lemma 4.8 proves that, with the aid of condition (28), this exponential decay persists even when $r=R$.
4.4.1. Monotonicity of $\nabla^{1}(\kappa,r)$
Let $v\in(\kappa)_{\iota}$ with $\iota\in\{I,II\}$ for $\kappa$ large enough. Note that if $v=v^{\prime}+v^{\prime\prime}$, then $\Theta_{1}(F^{j}(v))=\Theta_{1}(F^{j}(v^{\prime}))+\Theta_{1}(F^{j}(v^{\prime\prime}))$ for each $j\in\mathbb{Z}$. Moreover,
$\frac{\Theta_{1}(F^{j}(v))}{\Theta_{1}(v)}=\frac{\Theta_{1}(v^{\prime})}{\Theta_{1}(v)}\frac{\Theta_{1}(F^{j}(v^{\prime}))}{\Theta_{1}(v^{\prime})}+\frac{\Theta_{1}(v^{\prime\prime})}{\Theta_{1}(v)}\frac{\Theta_{1}(F^{j}(v^{\prime\prime}))}{\Theta_{1}(v^{\prime\prime})}.$
@@ 4.4.1. Monotonicity of $\nabla^{1}(\kappa,r)$
Lemma 4.4.
For any $r\leq\kappa$, we have
$\nabla^{1}(\kappa,r)\leq\nabla^{1}(\kappa+1,r)\leq\nabla^{1}(\kappa+1,r+1).$
Proof.
We recall that $s(v)=\frac{r(v)}{\Theta_{1}(v)}$. For any two chords $v$ and $v^{\prime}$, suppose $v$ can be decomposed as $v=v_{1}+v_{2}$. From the convex combination
$s(v_{1}+v_{2})=\frac{\Theta_{1}(v_{1})}{\Theta_{1}(v_{1})+\Theta_{1}(v_{2})}\,s(v_{1})+\frac{\Theta_{1}(v_{2})}{\Theta_{1}(v_{1})+\Theta_{1}(v_{2})}\,s(v_{2}),$
and by direct computation we obtain the decomposition of $\nabla^{1}(v_{1}+v_{2},v^{\prime})$:
$\nabla^{1}(v,v^{\prime})$ $=\frac{s(v^{\prime})-b_{1}s(v_{1})-b_{2}s(v_{2})}{\Theta(v_{1}+v_{2},v^{\prime})}$
$=\frac{\Theta_{2}(v_{1},v^{\prime})}{\Theta_{2}(v_{1}+v_{2},v^{\prime})}\,b_{1}\nabla^{1}(v_{1},v^{\prime})+\frac{\Theta_{2}(v_{2},v^{\prime})}{\Theta_{2}(v_{1}+v_{2},v^{\prime})}\,b_{2}\nabla^{1}(v_{2},v^{\prime})$
(35) $=c_{1}\nabla^{1}(v_{1},v^{\prime})+c_{2}\nabla^{1}(v_{2},v^{\prime}),$
where the coefficients are given by
$b_{1}=\frac{\Theta_{1}(v_{1})}{\Theta_{1}(v_{1})+\Theta_{1}(v_{2})},\qquad b_{2}=\frac{\Theta_{1}(v_{2})}{\Theta_{1}(v_{1})+\Theta_{1}(v_{2})},$
$c_{1}=\frac{\Theta_{2}(v_{1},v^{\prime})}{\Theta_{2}(v_{1}+v_{2},v^{\prime})}b_{1},\qquad c_{2}=\frac{\Theta_{2}(v_{2},v^{\prime})}{\Theta_{2}(v_{1}+v_{2},v^{\prime})}b_{2}.$
A direct calculation shows that $c_{1}+c_{2}=b_{1}+b_{2}=1$.
More generally, suppose $v$ and $v^{\prime}$ can be split as
$v=\sum_{i=1}^{I}v_{i},\qquad v^{\prime}=\sum_{j=1}^{J}v_{j}^{\prime},$
with $I,J\in\mathbb{N}$. Then
$\nabla^{1}(v,v^{\prime})=\sum_{i=1}^{I}d_{i}\Bigl(\sum_{j=1}^{J}e_{j}\nabla^{1}(v_{i},v_{j}^{\prime})\Bigr),$
i.e., $\nabla^{1}(v,v^{\prime})$ can be expressed as a convex combination of the quantities $\nabla^{1}(v_{i},v_{j}^{\prime})$.
Using the arithmetic properties of a rotation number of constant type, for given $v,v^{\prime}\in(\kappa,r)_{II}$ we can split $v$ and $v^{\prime}$ into several chords such that each $v_{i},v_{j}^{\prime}\in(\kappa+1,r)_{II}$. By definition we then obtain
$\nabla^{1}(\kappa,r)\leq\nabla^{1}(\kappa+1,r).$
By Lemma 4.2, if $(v_{1},v_{2})\in(\kappa+1,r)_{II}$, then
$\nabla^{1}(v_{1},v_{2})=\Theta_{2}(v_{1},v_{2})^{-1}\Bigl(\frac{\Theta_{1}(F(v_{2}))}{\Theta_{1}(v_{2})}-\frac{\Theta_{1}(F(v_{2}))}{\Theta_{1}(v_{2})}\Bigr).$
Again using the arithmetic properties of a constant-type rotation number, we can insert a chord $v_{3}\in(\kappa+1)_{II}$ between $v_{1}$ and $v_{2}$ so that
$(v_{1},v_{3})\in(\kappa+1,r+1)_{II},\qquad(v_{3},v_{2})\in(\kappa+1,r+1)_{II}.$
Note that
$\nabla^{1}(v_{1},v_{2})=\frac{\Theta_{2}(v_{1},v_{3})}{\Theta_{2}(v_{1},v_{2})}\,\nabla^{1}(v_{1},v_{3})+\frac{\Theta_{2}(v_{3},v_{2})}{\Theta_{2}(v_{1},v_{2})}\,\nabla^{1}(v_{3},v_{2}),$
$\Theta_{2}(v_{1},v_{2})=\Theta_{2}(v_{1},v_{3})+\Theta_{2}(v_{3},v_{2}).$
Hence, by definition,
$\nabla^{1}(\kappa+1,r)\leq\nabla^{1}(\kappa+1,r+1).$
4.4.2. Exponential decay of $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)$ with $r\ll R$
For simplicity of notation, we write
$F^{i}(v_{1},\dots,v_{4})\;:=\;\bigl(F^{i}(v_{1}),\dots,F^{i}(v_{4})\bigr),$
and we adopt analogous conventions for the action of $F$ on pairs of chords.
Denote
$K^{2}(v_{1},\ldots,v_{4}):=\frac{K^{1}(v_{3},v_{4})}{\Theta_{2}(v_{3},v_{4})}-\frac{K^{1}(v_{1},v_{2})}{\Theta_{2}(v_{1},v_{2})};$
$K^{2}(j|v_{1},\ldots,v_{4}):=\frac{K^{1}(j|v_{3},v_{4})}{\Theta_{2}(v_{3},v_{4})}-\frac{K^{1}(j|v_{1},v_{2})}{\Theta_{2}(v_{1},v_{2})}.$
It follows that
(36) $\Theta_{2}(v_{3},v_{4})K^{2}(j|v_{1},\ldots,v_{4})=K^{1}(j|v_{3},v_{4})-\frac{\Theta_{2}(v_{3},v_{4})}{\Theta_{2}(v_{1},v_{2})}K^{1}(j|v_{1},v_{2}).$
A direct calculation shows (see [11, Proposition 2.3])
Lemma 4.5.
Given $(v_{1},v_{2},v_{3},v_{4})\in(\kappa,r,s)_{II}$, we have
$K^{2}(v_{1},\ldots,v_{4})\lesssim\nabla^{2}(v_{1},\ldots,v_{4})+\Theta_{4}(v_{1},\ldots,v_{4}).$
Before proving the exponential decay of $\widetilde{\mathbf{K}}_{R}^{1}(\kappa,R)$, we first present a rough estimate that does not rely on condition (28). Building on this, we then use (28) to obtain the finer estimate in Lemma 4.3.
@@ 4.4.2. Exponential decay of $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)$ with $r\ll R$
    There holds $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}$.
@@
    $K^{2}(j|v_{1},\ldots,v_{4})-\sum_{i=0}^{j-1}K^{2}(F^{i}(v_{1},\ldots,v_{4}))\lesssim\varepsilon e^{\frac{\varepsilon}{10}R}q_{r}q_{R}^{-1}\sum_{i=0}^{j-1}\frac{\Theta_{2}(F^{i}(v_{3},v_{4}))}{\Theta_{2}(v_{3},v_{4})}.$
@@
    (37) $e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}\leq e^{\frac{\varepsilon}{10}R}q_{r}q_{R}^{-1}\sim e^{\frac{\varepsilon}{10}R}q_{R}^{-\frac{\varepsilon}{2}}<e^{\frac{\varepsilon}{10}R}q_{R}^{-\frac{\varepsilon}{3}}\leq(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
Proof.
By (11), $q_{R}\geq\sqrt{2}^{R}$. For Item (3), a direct calculation shows
@@
For $(v_{1},v_{2})\in(\kappa,r)_{II}$ with $0\leq r\leq R$, denote by $J(v_{1},v_{2})$ the arc on $\mathbb{T}$ spanned by $v_{1}\cup v_{2}$. By definition, we have
$\frac{1}{q_{r+2}}<\lambda_{2}(v_{1},v_{2})<\frac{1}{q_{r}}.$
Then for $0\leq j<q_{r}$, no point on $\mathbb{T}$ is contained in more than two of the arcs $\{F^{i}(J(v_{1},v_{2}))\}_{i=0}^{j}$ (see [7, p65, Proposition (8.3)]). It implies
(38) $\sum_{i=0}^{j}\Theta_{2}(F^{i}(v_{1},v_{2}))\leq 2.$
@@
$=\sum_{i=0}^{j-1}\left(\ln\frac{\Theta_{1}(F^{i+1}(v_{2}))}{\Theta_{1}(F^{i}(v_{2}))}-\ln\frac{\Theta_{1}(F^{i+1}(v_{1}))}{\Theta_{1}(F^{i}(v_{1}))}\right).$
By (16), $\frac{\Theta_{1}(F^{i+1}(v))}{\Theta_{1}(F^{i}(v))}$ is bounded for each $v\in(\kappa)_{II}$ and $i\in\mathbb{Z}$. It follows from the mean value theorem and Lemma 4.2 that for $0\leq j<q_{r}$,
(39) $|K^{1}(j|v_{1},v_{2})|\leq 3\nabla^{1}(\kappa,r)\sum_{i=0}^{j}\Theta_{2}(F^{i}(v_{1},v_{2}))\leq\frac{\varepsilon}{80A_{\alpha}}.$
Note that
(40) $r\leq R+2\gamma_{0}\leq\tilde{N}(\kappa)+2\gamma_{0}=n_{\kappa}.$
It follows from (27) and Lemma 4.4 that
$\widetilde{\mathbf{K}}_{r}^{1}(\phi(r),r)\leq\frac{\varepsilon}{80A_{\alpha}},\quad\widetilde{\mathbf{K}}_{r}^{1}(\kappa,r)\leq\frac{\varepsilon}{80A_{\alpha}},$
where $\phi$ is defined by (8). By Proposition 4.1 and (30), for $0<j<q_{R+2\gamma_{0}}$, where $\gamma_{0}$ is defined by (12), we have
$K^{0}(j|v)\leq 2a_{R+2\gamma_{0}}\widetilde{\mathbf{K}}_{R+2\gamma_{0}-1}^{1}(\kappa,R+2\gamma_{0}-1)+\ldots+2a_{l+1}\widetilde{\mathbf{K}}_{l}^{1}(\kappa,l)+\ldots+2a_{1}\widetilde{\mathbf{K}}_{0}^{1}(\kappa,0),$
where $1\leq a_{l}\leq A_{\alpha}$. Similar to the argument in (31), (32) and (33), we have
(41) $\widetilde{\mathbf{K}}_{R+2\gamma_{0}}^{0}(\kappa)\leq\frac{\varepsilon(R+c)}{40},$
where $0<c\leq q^{-1}_{R}q_{R+2\gamma_{0}}$. Due to the arithmetics of $\alpha$, $c$ is a constant independent of $R$.
By definition, for $(v_{1},v_{2})\in(\kappa,R)_{II}$, there exists a chord $\bar{v}$ determined by $v_{1},v_{2}$ such that $\lambda_{1}(\bar{v})=\lambda_{2}(v_{1},v_{2})$ and $\bar{v}\in({\phi(R)})_{I}$. Note from (34) that
$\widetilde{\mathbf{K}}_{R+2\gamma_{0}}^{0}({\phi(R)})\leq\widetilde{\mathbf{K}}_{R+2\gamma_{0}}^{0}(\kappa)\leq\frac{\varepsilon(R+c)}{40}.$
Since $\bar{v}\in({\phi(R)})_{I}$, then $\lambda_{1}(\bar{v})\geq\frac{2}{q_{R+2\gamma_{0}}}$. By Proposition 3.2, there exists $0\leq j_{0}<q_{R+2\gamma_{0}}$ such that
$\Theta_{1}(F^{j_{0}}(\bar{v}))\lesssim\lambda_{1}(\bar{v}).$
By the definition of $\widetilde{\mathbf{K}}_{R+2\gamma_{0}}^{0}({\phi(R)})$ (see (17) and (18)), for each $j$ with $|j-j_{0}|<q_{R+2\gamma_{0}}$, we have
$\frac{\varepsilon(R+c)}{40}\leq\ln\frac{\Theta_{1}(F^{j}(\bar{v}))}{\Theta_{1}(F^{j_{0}}(\bar{v}))}\leq\frac{\varepsilon(R+c)}{40}.$
Note that for each $i\in\mathbb{Z}$,
$F^{i}(\bar{v})=F^{i}(v_{1},v_{2}),\quad\lambda_{1}(\bar{v})=\lambda_{2}(F^{i}(v_{1},v_{2})).$
It follows that for each $j$ with $|j-j_{0}|<q_{R+2\gamma_{0}}$, we have
$\frac{\Theta_{2}(F^{j}(v_{1},v_{2}))}{\lambda_{2}(F^{j}(v_{1},v_{2}))}=\frac{\Theta_{1}(F^{j}(\bar{v}))}{\lambda_{1}(\bar{v})}\lesssim\frac{\Theta_{1}(F^{j}(\bar{v}))}{\Theta_{1}(F^{j_{0}}(\bar{v}))}\leq e^{\frac{\varepsilon(R+c)}{40}}.$
@@
$K^{1}(j|v_{1},v_{2})\lesssim\varepsilon\sum_{i=0}^{q_{r}}\Theta_{2}(F^{i}(v_{1},v_{2}))\leq\varepsilon e^{\frac{\varepsilon(R+c)}{40}}\sum_{i=0}^{q_{r}}\lambda_{2}(F^{i}(v_{1},v_{2}))\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}.$
@@
(42) $\widetilde{\mathbf{K}}_{r}^{1}(\phi(R),R)\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1},\quad\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}.$